Find the minimum sample size n needed to estimate mu for the given values of​ c, sigma​, and E. c= 0.98​, sigma= 8.7​, and E= 1

Find the margin of error for the given values of​ c, s, and n. c=0.98​, s=2.4​, n =25

Find the critical value Tc for the confidence level c=0.99 and sample size n=88.
Answer #1

Solution :

a ) Given that,

standard deviation = = 8.7

margin of error = E = 1

At 98% confidence level the z is ,

= 1 - 98% = 1 - 0.98 = 0.02

/ 2 = 0.02 / 2 = 0.01

Z/2 = Z0.01 = 2.326

Sample size = n = ((Z/2 * ) / E)2

= ((2.326 * 6250) / 1)2

= 410

Sample size = 410

b ) Given that,

= 3.2

s = 0.9

n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

At 98% confidence level the z is ,

= 1 - 98% = 1 - 0.98 = 0.02

/ 2 = 0.02 / 2 = 0.01

t /2,df = t0.01,24 =2.492

Margin of error = E = t/2,df * (s /n)

= 2.492* (2.4 / 25)

= 1.20

Margin of error = 1.20

c ) Given that,

n = 88

Degrees of freedom = df = n - 1 = 88 - 1 = 87

At 99% confidence level the t is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

t /2,df = t0.005,87 =2.633

The critcal value t =2.633